# What is my quantum computer good for? Quantum capability learning with physics-aware neural networks

### v1

@@ Abstract
Quantum computers have the potential to revolutionize diverse fields, including quantum chemistry, materials science, and machine learning. However, contemporary quantum computers experience errors that often cause quantum programs run on them to fail. Until quantum computers can reliably execute large quantum programs, stakeholders will need fast and reliable methods for assessing a quantum computer’s capability—i.e., the programs it can run and how well it can run them. Previously, off-the-shelf neural network architectures have been used to model quantum computers’ capabilities, but with limited success, because these networks fail to learn the complex quantum physics that determines real quantum computers’ errors. We address this shortcoming with a new quantum-physics-aware neural network architecture for learning capability models. Our architecture combines aspects of graph neural networks with efficient approximations to the physics of errors in quantum programs. This approach achieves up to $\sim 50\%$ reductions in mean absolute error on both experimental and simulated data, over state-of-the-art models based on convolutional neural networks.
@@ 1 Introduction
Quantum computers have the potential to efficiently solve classically intractable problems in quantum chemistry [Cao et al., 2019], materials science [Rubin et al., 2023], machine learning [Harrow et al., 2009], and cryptography [Shor, 1997]. While contemporary quantum computers are approaching the size and noise levels needed to solve interesting problems [Arute et al., 2019], they are far from being capable of reliably running most useful quantum programs [Proctor et al., 2021a]. Until we build quantum computers capable of executing any and all useful and interesting quantum programs, stakeholders will require fast, reliable, and scalable methods for predicting the programs that a given quantum computer can reliably execute.
@@
The task of learning which quantum programs a particular quantum computer can reliably execute is known as quantum capability learning [Proctor et al., 2021a]. Quantum capability learning is very difficult because the number of possible (Markovian) errors plaguing a quantum computer grows exponentially in its size [Blume-Kohout et al., 2022], i.e., in the number of qubits ($n$) it contains, and errors in a quantum program can combine in difficult-to-predict ways [Proctor et al., 2021a]. Most existing approaches to capability learning restrict themselves to learning how well a quantum computer executes a small set of quantum programs, by running all of those programs and estimating a success metric for each one [Lubinski et al., 2023]. While these methods provide insight into a quantum computer’s capability, they are not predictive.
@@
Recently, several groups have proposed building predictive models of a quantum computer’s capability using convolutional neural networks (CNNs) [Elsayed Amer et al., 2022, Hothem et al., 2023c, Vadali et al., 2024, Hothem et al., 2023b] and graph neural networks (GNNs) [Wang et al., 2022]. However, these neural-network-based capability models achieve only modest prediction accuracy when applied to real quantum computers, because they fail to learn the complex physics that determines real quantum computers’ failures [Hothem et al., 2023c].
@@
In this work, we introduce a novel quantum-physics-aware neural network (qpa-NN) architecture for quantum capability learning (Fig. 1). Our approach uses neural networks with GNN-inspired structures to predict the rates of the most physically relevant errors in quantum programs. These predicted error rates are then combined using an efficient approximation to the exact (but exponentially costly) quantum physics formula for how those errors combine to impact a program’s success rate. Our approach leverages the graph structures that encode the physics of how errors’ rates typically depend on both the quantum program being run and how a quantum computer’s qubits are arranged, and it offloads the difficult-to-learn, yet classically tractable task of approximately combining these error rates to predict a circuit’s performance to an already-known function. This enables our qpa-NNs to vastly outperform the state-of-the-art CNNs of Hothem et al. [2023c] on both experimental and simulated data.
@@
In a head-to-head comparison, our qpa-NNs achieve a $\sim 50\%$ reduction in mean absolute error (MAE) over the CNNs of Hothem et al. [2023c], on average and on the same experimental datasets. Our qpa-NNs achieve an average $\sim 36\%$ improvement over those CNNs even after fine-tuning those CNNs on the same subset of the training data (high-fidelity programs) used to train our qpa-NNs.
Our qpa-NNs’ improved performance is likely largely due to their improved ability to model the impact of coherent errors on a program’s success rate. Off-the shelf networks struggle with coherent errors [Hothem et al., 2023c], but our qpa-NNs are designed to model how these errors add up and cancel out, making the qpa-NNs much better predictors in the presence of coherent errors. To verify this, we demonstrate that our qpa-NNs can accurately predict the performance of random circuits run on a hypothetical 4-qubit quantum computer experiencing only coherent errors. Our qpa-NN obtained a $\sim 50\%$ lower MAE than a CNN, averaged across five datasets, and the trained qpa-NN even exhibits moderate performance when making predictions for a different class of circuits (random mirror circuits [Proctor et al., 2021a]) simulated on the same hypothetical 4-qubit quantum computer, i.e., our qpa-NNs display moderate prediction accuracy on out-of-distribution data.
@@ 2 Background
In this section, we review the background in quantum computing necessary to understand this paper. See Nielsen and Chuang [2010] for an in-depth introduction to quantum computing and Blume-Kohout et al. [2022] for a thorough description of the errors in quantum computers.
@@ 2.2 Quantum capability learning
Because quantum computers are error-prone, knowing which quantum circuits a particular quantum computer can execute with low error probability is important. Known as quantum capability learning [Proctor et al., 2021a, Hothem et al., 2023c], this task formally involves learning the mapping between a set of quantum circuits $c\in\mathcal{C}$ and some success metric $s(c)\in\mathbb{R}$ quantifying how well $c$ runs on a quantum computer $\mathcal{Q}$. In this work we consider a large class of circuits known as Clifford (or stabilizer) circuits [Aaronson and Gottesman, 2004], which are sufficient to enable quantum error correction [Campbell et al., 2017], and two widely-used success metrics: probability of successful trial (PST) (a.k.a., success probability) and the process fidelity (a.k.a. entanglement fidelity) [Hothem et al., 2023c, Nielsen, 2002].
@@
Estimating $F(c)$ is more complicated than estimating $\mathrm{PST}(c)$ (it requires embedding $c$ into other circuits), but efficient methods exist [Proctor et al., 2022].
@@ 2.3 Modelling errors in quantum computers
Equations (4) and (5) are good approximations for low-error circuits [Mądzik et al., 2022].
@@ 3.1 Physics-aware neural networks for predicting errors in quantum circuits
The internal structures of $\mathcal{N}$ are chosen to reflect general physical principles for how $\mathcal{E}$ and $\vec{m}$ depend on $c$. $\mathcal{E}_{ij}$ is a prediction for the rate that $G_{j}$ occurs in circuit layer $i$, and this error corresponds to a space/time location within $c$—because it occurs at layer index or time $i$ and $G_{j}$ acts on a subset of the qubits $Q(G_{j})$ (see example in Fig. 1c). This error’s rate will therefore primarily depend only on the gates in a time- and space-local region around its location in $c$. Furthermore this dependence will typically be invariant under time translations (this is true except for some exotic non-Markovian kinds of errors, that we discuss in Section 7.1). We can encode these structures into $\mathcal{N}$ by predicting $\mathcal{E}_{ij}$ from a space-time “window” of $c$ around the associated error’s location using a filter $W_{j}$ that “slides” across the circuit to predict the rate of $G_{j}$ versus time $i$. Stated more formally, we predict $\mathcal{E}_{ij}$ using a multilayer perceptron $\mathcal{N}_{j}$ whereby $\mathcal{N}_{j}(W_{j}[I(c),i])=\mathcal{E}_{ij}$ and $W_{j}[I(c),i]$ is a snippet of $I(c)$ whose temporal origin is $i$ (see Fig. 1e). The shape of each filter $W_{j}$ is a hyperparameter of our networks and it can be designed to reflect general physical principles, the known physics of a particular quantum computing system, and/or optimized with hyperparameter tuning. The particular neural networks we present later herein use filters $W_{j}(I(c),i)$ that snip out only layer $i$ and discard the parts of the layer that act on qubits more than $l$ steps away from $Q(G_{j})$ in the quantum computer’s connectivity graph (e.g., the filter shown in Fig. 1e corresponds to the error shown in Fig. 1c and $l=1$). This neural network structure has close connections to graph convolution layers [Kipf and Welling, 2016], as well as CNNs. We choose this structure as it can model spatially localized crosstalk errors, which are a ubiquitous but hard-to-model class of errors in quantum computers [Sarovar et al., 2020].
@@
The network $\mathcal{N}$ must also predict the rates of errors that occur during measurements (unless the qpa-NN will only ever predict $F(c)$ not $\mathrm{PST}(c)$), but these are typically independent of the rates of gate errors (which are predicted by the $\mathcal{N}_{j}$). So we do not use the $\mathcal{N}_{j}$ and their convolutional filters $W_{j}$ to make predictions for $\vec{m}$. Instead we use separate but structurally equivalent networks $\mathcal{N}^{\prime}_{j}$ with corresponding filters $W^{\prime}_{j}$ that take $M(c)$ as input and implement only spatial filtering. That is, $W^{\prime}_{j}$ simply discard rows from $M(c)$, as, unlike $I(c)$, $M(c)$ has no temporal dimension. The $W_{j}^{\prime}$ are hyperparameters of our networks and we can separately adjust the shape of each $W^{\prime}_{j}$ to reflect the known physics of errors induced by measuring qubits. In our demonstrations, our $W_{j}^{\prime}$ filters have the same structure as the $W_{j}$ filters but with an independent $l^{\prime}$ steps parameter (large $l^{\prime}$ enables modelling many-qubit measurement crosstalk).
@@ 3.2 Processing predicted error rates to predict capabilities
We process $\mathcal{N}$’s output to predict $\mathrm{PST}(c)$ or $F(c)$ using a function $f$ with no learnable parameters. This turns $\mathcal{N}$’s output into the two quantities of interest, and it is also makes training $\mathcal{N}$ feasible. We cannot easily train $\mathcal{N}$ in isolation because the error matrix $\mathcal{E}$ predicted by $\mathcal{N}$ is not a directly observable quantity. Generating the data needed to train $\mathcal{N}$ directly would require extraordinarily expensive quantum process tomography [Nielsen et al., 2021], which is infeasible except for very small $n$. In contrast, both $\mathrm{PST}(c)$ and $F(c)$ can be efficiently estimated (see Section 2) for a given circuit $c$.
@@ 4.1 Experimental data
We used the 5-qubit datasets in Hothem et al. [2023c] for our experimental demonstrations. Each of these datasets $D=\{(c,\widehat{\mathrm{PST}(c)})\}$ was gathered by running random and periodic mirror circuits (two types of definite-outcome circuits) on 5-qubit IBM Q computers ($\mathtt{ibmq\_london}$, $\mathtt{ibmq\_essex}$, $\mathtt{ibmq\_burlington}$, $\mathtt{ibmq\_vigo}$, $\mathtt{ibmq\_ourense}$ and $\mathtt{ibmq\_yorktown}$), and estimating the PST of each circuit using Eq. (2). Each circuit was run between $1024$ and $4096$ times, with the exact number depending upon how many times the circuit sampling process generated the circuit (some short, $1$-qubit circuits were generated multiple times). The random and periodic mirror circuits contained between $1$ and $5$ active qubits—called the circuit’s width—and ranged in depth from $3$ to $515$ layers (alt. $259$ for the $\mathtt{ibmq\_yorktown}$ dataset).
@@
In this work we focus on high-PST circuits, so we removed all circuits with a PST less than $85\%$ from each dataset, leaving between $864$ ($\mathtt{ibmq\_burlington}$) and 1369 ($\mathtt{ibmq\_yorktown}$) circuits in each dataset. The remaining circuits were partitioned into training, validation, and test sets based upon their original assignment in Hothem et al. [2023c]. This setup enables a direct comparison between our qpa-NNs and the CNNs trained in Hothem et al. [2023c]. Training set sizes ranged from $682$ circuits on $\mathtt{ibmq\_burlington}$ to $1097$ circuits on $\mathtt{ibmq\_yorktown}$, with an approximate training, validation, testing split of $80\%$, $10\%$, and $10\%$, respectively.
4.2 Simulated data
@@
For our simulations, we generated 5 datasets of $5000$ high-fidelity ($F(c)>85\%$) random circuits, for a hypothetical 4-qubit processor with a “ring” geometry (i.e., like that in Fig. 1b). The circuits ranged in width ($w$) from $1$ to $4$ qubits, and in depth from 1 to 180 circuit layers. Each circuit was designed for a randomly chosen subset of $w$ qubits. Each circuit layer was created by i.i.d. sampling from all possible circuit layers on the $w$ active qubits. We used a gate set containing two-qubit $\mathrm{CNOT}$ gates and 7 different single-qubit gates (specifically $\{X(\pi/2),Y(\pi/2),X(3\pi/2),Y(3\pi/2),X(\pi),Y(\pi),Z(\pi)\}$ where $P(\theta)$ denotes a rotation around the $P$ axis by $\theta$). See Appendix C for additional details.
@@
All circuits were simulated under the same error model, consisting of local coherent (i.e., H) errors, to exactly compute each $c$’s $F(c)$ [Fig. 3 shows a histogram of $F(c)$]. After removing duplicate circuits, the resulting datasets $D=\{(c,F(c))\}$ were partitioned into training, validation, and testing subsets, with a partition of $56.25\%$, $18.75\%$, and $25\%$, respectively. The parameters of the error model were randomly selected: each gate was assigned a small error strength, which was then distributed randomly across all possible (local) one- or two-qubit coherent errors, for the one- and two-qubit gate, respectively. We chose a model with only coherent errors as these errors are ubiquitous, they are hard to model accurately and efficiently, and we conjecture that qpa-NNs can model them.
@@
We also generated 5 datasets of 750 random mirror circuits on the same hypothetical quantum computer. Again, random mirror circuits varied in width from $1$ to $4$ qubits, and were designed to be run on a randomly selected subset of $w$ qubits. However, instead of i.i.d. sampling of each circuit layer, each circuit was randomly sampled from the class of random mirror circuits on the $w$ qubits. The depth of the mirror circuits ranged from 8 to 174 layers. We generated the mirror circuit datasets to evaluate how well qpa-NNs and CNNs generalize to out-of-distribution circuits, so they were used exclusively as testing sets. To ensure that no training was performed on mirror circuits, we removed any mirror circuits that appeared in the random circuit sets (in actuality, there were no duplicates).

### v2

@@ Abstract
Quantum computers have the potential to revolutionize diverse fields, including quantum chemistry, materials science, and machine learning. However, contemporary quantum computers experience errors that often cause quantum programs run on them to fail. Until quantum computers can reliably execute large quantum programs, stakeholders will need fast and reliable methods for assessing a quantum computer’s capability—i.e., the programs it can run and how well it can run them. Previously, off-the-shelf neural network architectures have been used to model quantum computers’ capabilities, but with limited success, because these networks fail to learn the complex quantum physics that determines real quantum computers’ errors. We address this shortcoming with a new quantum-physics-aware neural network architecture for learning capability models. Our scalable architecture combines aspects of graph neural networks with efficient approximations to the physics of errors in quantum programs. This approach achieves up to $\sim 50\%$ reductions in mean absolute error on both experimental and simulated data, over state-of-the-art models based on convolutional neural networks, and scales to devices with 100+ qubits.
@@ 1 Introduction
Quantum computers have the potential to efficiently solve classically intractable problems in quantum chemistry [Cao et al., 2019], materials science [Rubin et al., 2024], machine learning [Harrow et al., 2009], and cryptography [Shor, 1997]. While contemporary quantum computers are approaching the size and noise levels needed to solve interesting problems [Arute et al., 2019], they are far from being capable of reliably running most useful quantum programs [Proctor et al., 2021a]. Until we build quantum computers capable of executing any and all useful and interesting quantum programs, stakeholders will require fast, reliable, and scalable methods for predicting the programs that a given quantum computer can reliably execute.
@@
The task of learning which quantum programs a particular quantum computer can reliably execute is known as quantum capability learning [Proctor et al., 2021a]. Quantum capability learning is very difficult because the number of possible (Markovian) errors plaguing a quantum computer grows exponentially in its size [Blume-Kohout et al., 2022], i.e., in the number of qubits ($n$) it contains, and errors in a quantum program can combine in difficult-to-predict ways [Proctor et al., 2021a]. Most existing approaches to capability learning restrict themselves to learning how well a quantum computer executes a small set of quantum programs, by running all of those programs and estimating a success metric for each one [Lubinski et al., 2023, Proctor et al., 2024]. While these methods provide insight into a quantum computer’s capability, they are not predictive.
@@
Recently, several groups have proposed building predictive models of a quantum computer’s capability using convolutional neural networks (CNNs) [Elsayed Amer et al., 2022, Hothem et al., 2024b, Vadali et al., 2024, Hothem et al., 2023b] and graph neural networks (GNNs) [Wang et al., 2022]. However, these neural-network-based capability models achieve only modest prediction accuracy when applied to real quantum computers, because they fail to learn the complex physics that determines real quantum computers’ failures [Hothem et al., 2024b].
@@
In this work, we introduce a novel quantum-physics-aware neural network (qpa-NN) architecture for quantum capability learning (Fig. 1). Our approach uses neural networks with GNN-inspired structures to predict the rates of the most physically relevant errors in quantum programs. These predicted error rates are then combined using an efficient approximation to the exact (but exponentially costly) quantum physics formula for how those errors combine to impact a program’s success rate. Our approach leverages the graph structures that encode the physics of how errors’ rates typically depend on both the quantum program being run and how a quantum computer’s qubits are arranged, and it offloads the difficult-to-learn, yet classically tractable task of approximately combining these error rates to predict a circuit’s performance to an already-known function. This enables our qpa-NNs to vastly outperform the state-of-the-art CNNs of Hothem et al. [2024b] on both experimental and simulated data without sacrificing the ability to model large devices of 100+ qubits.
@@
In a head-to-head comparison, our qpa-NNs achieve a $\sim 50\%$ reduction in mean absolute error (MAE) over the CNNs of Hothem et al. [2024b], on average and on the same experimental datasets. Our qpa-NNs achieve an average $\sim 36\%$ improvement over those CNNs even after fine-tuning those CNNs on the same subset of the training data (high-fidelity programs) used to train our qpa-NNs.
Our qpa-NNs’ improved performance is likely largely due to their improved ability to model the impact of coherent errors on a program’s success rate. Off-the-shelf networks struggle with coherent errors [Hothem et al., 2024b], but our qpa-NNs are designed to model how these errors add up and cancel out, making the qpa-NNs much better predictors in the presence of coherent errors. To verify this, we demonstrate that our qpa-NNs can accurately predict the performance of random circuits run on a hypothetical 4-qubit quantum computer experiencing only coherent errors. Our qpa-NN obtained a $\sim 50\%$ lower MAE than a CNN, averaged across five datasets, and the trained qpa-NN even exhibits moderate performance when making predictions for a different class of circuits (random mirror circuits [Proctor et al., 2021a]) simulated on the same hypothetical 4-qubit quantum computer, i.e., our qpa-NNs display moderate prediction accuracy on out-of-distribution data.
We make the following contributions in our work:
  1. 1.
    We introduce qpa-NNs, a bespoke neural network architecture for modeling the capability of a quantum computer, which outperform state-of-the-art CNN models by $\sim 50\%$ on experimental and simulated data.
  2. 2.
    We use our qpa-NNs to model the capability of a simulated 100-qubit device; the largest ever neural network capability learning demonstration by a factor of two.
  3. 3.
    We demonstrate, for the first time, how to train NNs to predict the process fidelity [Nielsen, 2002] of a circuit, which is the most widely used quantum channel error metric.
  4. 4.
    We provide evidence that the improved performance of our qpa-NNs is partly due to their ability to better model the effect of coherent errors, which are known to be challenging for other state-of-the-art methods.
@@ 2 Background
In this section, we review the background in quantum computing necessary to understand this paper. See Nielsen and Chuang [2010] for an in-depth introduction to quantum computing and Blume-Kohout et al. [2022] or Hashim et al. [2024] for a thorough description of the errors in quantum computers.
@@ 2.2 Quantum capability learning
Because quantum computers are error-prone, knowing which quantum circuits a particular quantum computer can execute with low error probability is important. Known as quantum capability learning [Proctor et al., 2021a, Hothem et al., 2024b], this task formally involves learning the mapping between a set of quantum circuits $c\in\mathcal{C}$ and some success metric $s(c)\in\mathbb{R}$ quantifying how well $c$ runs on a quantum computer $\mathcal{Q}$. In this work, we consider a large class of circuits known as Clifford (or stabilizer) circuits [Aaronson and Gottesman, 2004], which are sufficient to enable quantum error correction [Campbell et al., 2017], and two widely-used success metrics: probability of successful trial (PST) (a.k.a. success probability) and the process fidelity (a.k.a. entanglement fidelity) [Hothem et al., 2024b, Nielsen, 2002].
@@
Estimating $F(c)$ is more complicated than estimating $\mathrm{PST}(c)$, but efficient methods exist, such as mirror circuit fidelity estimation [Proctor et al., 2022]. Hence, in theory, it is possible to efficiently gather training data using either $\mathrm{PST}(c)$ or $F(c)$ on arbitrarily large quantum computers.
@@ 2.3 Modelling errors in quantum computers
Equations (4) and (5) are good approximations for low-error circuits [Mądzik et al., 2022]. However, they both suffer from the same flaw: they require tracking $\mathcal{O}(4^{n})$ parameters. To address this problem, our qpa-NNs make an approximation: they only account for the contributions of a polynomially-sized set of errors that contains all those errors which are most likely to be experienced by a quantum computer. In this work, we chose to account for only local, low-weight errors, i.e., those with initial support on a small, connected subset of a device’s connectivity graph.
@@ 3.1 Physics-aware neural networks for predicting errors in quantum circuits
The internal structures of $\mathcal{N}$ are chosen to reflect general physical principles for how $\mathcal{E}$ and $\vec{m}$ depend on $c$. $\mathcal{E}_{ij}$ is a prediction for the rate that $G_{j}$ occurs in circuit layer $i$, and this error corresponds to a space/time location within $c$—because it occurs at layer index or time $i$ and $G_{j}$ acts on a subset of the qubits $Q(G_{j})$ (see example in Fig. 1c). This error’s rate will therefore primarily depend only on the gates in a time- and space-local region around its location in $c$. Furthermore, this dependence will typically be invariant under time translations (this is true except for some exotic non-Markovian kinds of errors, which we discuss in Section 8.1). We can encode these structures into $\mathcal{N}$ by predicting $\mathcal{E}_{ij}$ from a space-time “window” of $c$ around the associated error’s location using a filter $W_{j}$ that “slides” across the circuit to predict the rate of $G_{j}$ versus time $i$. Stated more formally, we predict $\mathcal{E}_{ij}$ using a multilayer perceptron $\mathcal{N}_{j}$ whereby $\mathcal{N}_{j}(W_{j}[I(c),i])=\mathcal{E}_{ij}$ and $W_{j}[I(c),i]$ is a snippet of $I(c)$ whose temporal origin is $i$ (see Fig. 1e). The shape of each filter $W_{j}$ is a hyperparameter of our networks and it can be designed to reflect general physical principles, the known physics of a particular quantum computing system, and/or optimized with hyperparameter tuning. The particular neural networks we present later herein use filters $W_{j}(I(c),i)$ that snip out only layer $i$ and discard the parts of the layer that act on qubits more than $l$ steps away from $Q(G_{j})$ in the quantum computer’s connectivity graph (e.g., the filter shown in Fig. 1e corresponds to the error shown in Fig. 1c and $l=1$). This neural network structure has close connections to graph convolution layers [Kipf and Welling, 2016], as well as CNNs. We choose this structure as it can model spatially localized crosstalk errors, which are a ubiquitous but hard-to-model class of errors in quantum computers [Sarovar et al., 2020].
@@
The network $\mathcal{N}$ must also predict the rates of errors that occur during measurements (unless the qpa-NN will only ever predict $F(c)$ not $\mathrm{PST}(c)$), but these are typically independent of the rates of gate errors (which are predicted by the $\mathcal{N}_{j}$). So we do not use the $\mathcal{N}_{j}$ and their convolutional filters $W_{j}$ to make predictions for $\vec{m}$. Instead we use separate but structurally equivalent networks $\mathcal{N}^{\prime}_{j}$ with corresponding filters $W^{\prime}_{j}$ that take $M(c)$ as input and implement only spatial filtering. That is, $W^{\prime}_{j}$ simply discards rows from $M(c)$, as, unlike $I(c)$, $M(c)$ has no temporal dimension. The $W_{j}^{\prime}$ are hyperparameters of our networks allowing us to separately adjust the shape of each $W^{\prime}_{j}$ to reflect the known physics of errors induced by measuring qubits. In our demonstrations, our $W_{j}^{\prime}$ filters have the same structure as the $W_{j}$ filters but with an independent $l^{\prime}$ steps parameter (large $l^{\prime}$ enables modelling many-qubit measurement crosstalk).
@@ 3.2 Processing predicted error rates to predict capabilities
We process $\mathcal{N}$’s output to predict $\mathrm{PST}(c)$ or $F(c)$ using a function $f$ with no learnable parameters. This turns $\mathcal{N}$’s output into the two quantities of interest, and it also makes training $\mathcal{N}$ feasible. We cannot easily train $\mathcal{N}$ in isolation because the error matrix $\mathcal{E}$ predicted by $\mathcal{N}$ is not a directly observable quantity. Generating the data needed to train $\mathcal{N}$ directly would require extraordinarily expensive quantum process tomography [Nielsen et al., 2021], which is infeasible except for very small $n$. In contrast, both $\mathrm{PST}(c)$ and $F(c)$ can be efficiently estimated (see Section 2) for a given circuit $c$.
@@ 4.1 Experimental 5-qubit data
We used the 5-qubit datasets from Hothem et al. [2024b] for our experimental demonstrations. Each of these datasets $D=\{(c,\widehat{\mathrm{PST}(c)})\}$ was gathered by running random and periodic mirror circuits (two types of definite-outcome circuits) on 5-qubit IBM Q computers ($\mathtt{ibmq\_london}$, $\mathtt{ibmq\_essex}$, $\mathtt{ibmq\_burlington}$, $\mathtt{ibmq\_vigo}$, $\mathtt{ibmq\_ourense}$ and $\mathtt{ibmq\_yorktown}$), and estimating the PST of each circuit using Eq. (2). Each circuit was run between $1024$ and $4096$ times, with the exact number depending upon how many times the circuit sampling process generated the circuit (some short, $1$-qubit circuits were generated multiple times). The random and periodic mirror circuits contained between $1$ and $5$ active qubits—called the circuit’s width—and ranged in depth from $3$ to $515$ layers (alt. $259$ layers for the $\mathtt{ibmq\_yorktown}$ dataset).
@@
As we focus on high-PST circuits, we removed all circuits with a PST less than $85\%$ from each dataset, leaving between $864$ ($\mathtt{ibmq\_burlington}$) and $1369$ ($\mathtt{ibmq\_yorktown}$) circuits in each dataset. The remaining circuits were partitioned into training, validation, and test sets by their original assignment in Hothem et al. [2024b]. This setup enables a direct comparison between our qpa-NNs and the CNNs trained in Hothem et al. [2024b]. Training set sizes ranged from $682$ circuits on $\mathtt{ibmq\_burlington}$ to $1097$ circuits on $\mathtt{ibmq\_yorktown}$, with an approximate training, validation, testing split of $80\%$, $10\%$, and $10\%$, respectively.
4.2 Simulated 4-qubit data
@@
For our 4-qubit simulations, we generated 5 datasets of $5000$ high-fidelity ($F(c)>85\%$) random circuits, for a hypothetical 4-qubit processor with a “ring” geometry (i.e., like that in Fig. 1b). The circuits ranged in width ($w$) from $1$ to $4$ qubits, and in depth from $1$ to $180$ circuit layers. We designed each circuit for a randomly chosen subset of $w$ qubits. Each circuit layer was created by i.i.d. sampling from all possible circuit layers on the $w$ active qubits. We used a gate set containing two-qubit $\mathrm{CNOT}$ gates and 7 different single-qubit gates (specifically $\{X(\pi/2),Y(\pi/2),X(3\pi/2),Y(3\pi/2),X(\pi),Y(\pi),Z(\pi)\}$ where $P(\theta)$ denotes a rotation around the $P$ axis of the Bloch sphere by $\theta$). See Appendix C for additional details.
@@
All circuits were simulated under the same error model, consisting of local coherent (i.e., H) errors, to exactly compute each $c$’s $F(c)$ [Fig. 3 shows a histogram of $F(c)$]. After removing duplicate circuits, the resulting datasets $D=\{(c,F(c))\}$ were partitioned into training, validation, and testing subsets, with a partition of $56.25\%$, $18.75\%$, and $25\%$, respectively. The parameters of the error model were randomly selected: each gate was assigned a small error strength, which was then distributed randomly across all possible (local) one- or two-qubit coherent errors, for the one- and two-qubit gates, respectively. We chose a model with only coherent errors as these errors are ubiquitous, they are hard to model accurately and efficiently, and we conjecture that qpa-NNs can model them.
@@
We also generated $5$ datasets of $750$ random mirror circuits on the same hypothetical 4-qubit quantum computer. Again, the random mirror circuits varied in width from $1$ to $4$ qubits, and were designed to be run on a randomly selected subset of $w$ qubits. However, instead of i.i.d. sampling of each circuit layer, each circuit was randomly sampled from the class of random mirror circuits on the $w$ qubits. The depth of the mirror circuits ranged from $8$ to $174$ layers. Because we generated the mirror circuit datasets to evaluate how well qpa-NNs and CNNs generalize to out-of-distribution circuits, they were used exclusively as testing sets. To ensure that no training was performed on mirror circuits, we removed any mirror circuits that appeared in the random circuit sets (in actuality, there were no duplicates).
4.3 Simulated 100-qubit data
For our 100-qubit simulation, we generated a single dataset of $5000$ high-fidelity ($F(c)>91\%$) random circuits, for a hypothetical 100-qubit quantum computer with a “ring” geometry. All of the circuits had a width of 100 qubits, and ranged in depth from 1 to 22 circuit layers. We sampled circuit layers using the same process and gate set as in the 4-qubit simulations.
We simulated every circuit using the same error model, consisting of local, weight-1 S and H errors. As before, the parameters of the error model were randomly selected and the data were partitioned into training, validation, and testing subsets according to a $56.25\%$, $18.75\%$, $25\%$ split.
Unlike in our 4-qubit simulations, we did not compute $F(c)$ exactly as doing so for a 100-qubit circuit is infeasible in the presence of coherent errors. Instead, we used a first-order simulation method to approximate $F(c)$. In this method, $F(c)$ is computed by assigning each gate its own error vector based on the error model, adding up the error vectors layer-wise to compute an error vector for each circuit layer, and then computing $F(c)$ as in the second part of a qpa-NN [Figure 1(f)]. See Appendix C.4 for more details.